Frankie and Johnny begin reading a novel at the same time. Frankie reads 8 pages a day and Johnny 5 pages a day. If Frankie is on page 72, what page is Johnny on?
Answer by checkley77(12844)

72/8=x/5
8x=72*5
8x=360
x=360/8
x=45 is the page Johnny is on.
Proof:
72/8=45/5
9=9